OpenStudy (ronlover101):

i need help how to find the x and y intercepts for these slope. Here are the questions: 1) x+2y=8 2) 2x + 2y +-4 3) 3x + y = 12 4) 2x + 4y=-4

OpenStudy (anonymous):

So first we need to know exactly what an x and y intercept is. An x intercept is the value of x when y = 0. This makes sense because it touches the x axis at this point. So with this we can look at the first problem...

OpenStudy (anonymous):

\[x + 2y = 8\] Let's set y=0 and solve for x to find the x intercept. \[x + 2(0) = 8\] So x = 8. if you have a graphing calculator you can check this too! ------------------------ For y intercept we set x=0. \[x + 2y = 8\] \[0 + 2y = 8\] \[y = 4\]

OpenStudy (ronlover101):

Thank you

OpenStudy (anonymous):

No problem, I'll be around here should you have any trouble!

OpenStudy (ronlover101):